OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

4x -y -{4*(-3) - 6}=0

OpenStudy (anonymous):

are you sure that is correct? the answer you came up with was not one of the choices on my hw.. please explain i need help

OpenStudy (anonymous):

4x-y +18=0

OpenStudy (anonymous):

write down the options...

OpenStudy (anonymous):

4x-y=-18

OpenStudy (anonymous):

x-4y=-18 x+4y=18 4x-y=-18 4x+4y=18

OpenStudy (anonymous):

3rd one...

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

:)